the force between two 2.11 x 10^-4 C charges that are 2.11 meters apart is ______N?
Answer #1

We know that force between the two charge

F=kq1q2/r

Here q1=q2=2.11*10-4 C , r=2.11 m , k=9*109 Nm2/C2

F=9*109*2.11*10-4*2.11*10-4/2.112

F=9*101

F=90 N